OpenStudy (anonymous):

find the indefinite integral of (5x + 21)/(x^2 + 10x + 25) dx

OpenStudy (turingtest):

are you familiar with partial fractions?

OpenStudy (anonymous):

not at all, i didn't figure i would have to use it for this question

OpenStudy (anonymous):

i evaluated it earlier using integration by parts and got \[5\ln (x + 5) - \frac{ 5x + 21 }{ x + 5 }\] it differentiates back to the original function but it's not what i get on my TI...

OpenStudy (turingtest):

it's likely that you can manipulate the answer your TI gives you into the same, could you type it please?

OpenStudy (anonymous):

sure, my TI gives me \[5\ln(x+5) + \frac{ 4 }{ x+5 }\]

OpenStudy (phi):

they look different. I would go with the TI on this one.

OpenStudy (turingtest):

I'd agree, you likely made a mistake

OpenStudy (anonymous):

hmm, i suppose so can this function not be integrated without using partial fractions? maybe that's where i messed up

OpenStudy (turingtest):

I think pf is really the way to go here

OpenStudy (phi):

This is surprising, but if we ask is there a constant (of integration) that makes these 2 expression equal: \[ - \frac{ 5x + 21 }{ x + 5 } = \frac{ 4 }{ x + 5 } + C\] we find \[ C= \frac{ -5x -21 - 4 }{ x + 5 } = \frac{ -5(x+5)}{(x+5)} = -5 \] for x≠5 so the two results are equal (to a constant)

OpenStudy (anonymous):

i was on the verge of that revelation myself just now, but in much more abstract terms--thanks for clearing that up